OpenStudy (anonymous):

how to factor 7t(n-4)-(4-n)

OpenStudy (anonymous):

I would rewrite the 2nd portion 7t (n-4) + (n-4) = 0 now combine like terms (7t + 1) (n-4) = 0

OpenStudy (anonymous):

how did you do that?

OpenStudy (anonymous):

Like I said, I would rewrite the 2nd portion. 4-n multiply each term by -1 = n -4 so I then "pull that -1" out 7t(n-4)-(4-n) = 0 7T (N-4) -(-1) (N-4) = 0 7T (N-4) +1 (N-4) = 0 what is common? so factor it out (7T +1) (N-4) = 0 think of it as distribution in reverse

OpenStudy (jdoe0001):

-(4-n) => -1*4 -1*-n => -4 +n => (n-4)

OpenStudy (jdoe0001):

recall same signs multiplied give + different signs multiplied give -